OpenStudy (anonymous):

prediction equation of a line for the ordered pair (4,4878) and (11,5616) ? prediction equation of a line for the ordered pair (4,4878) and (11,5616) ? @Mathematics

OpenStudy (anonymous):

Slope of the line is y2-y1/x2-x1, so 5616 - 4878/11 - 4 = 738/7 = m = slope. So, plug that slope, along with one of the coordinates (I'll use (4,4878)) in the point-slope formula \[y-y _{1}=m(x-x _{1})\] to get y-4878=738/7(x-4) which equals y-4878=738x/4 - 2952/11. Isolate y to get: \[y=738x/4 + 56610/11\] This equation look sloppy but it works. You could also of plotted the points on a scatterplot and used linear regression to find the same line.